Specific volume 7.40 m3 of air per minute at 31˚C DBT and 18.5˚C WBT is passed over the cooling

coil whose surface temperature is 4.4˚C. The coil cooling capacity is 3.56 tons of refrigeration

under the given condition of air. Determine DBT and WBT of the air leaving the cooling coil.


1
Expert's answer
2021-05-25T02:12:02-0400

Mass flow rate = 7.40 m3 /min = 1.23 *7.40 =9.102 kg/min = 0.1517 kg/s

Heat observed by the coil from air

"\bar{q}= \frac{12.5}{0.1517}=-82.39947" The negative sign shows heat is absorbed.

"q=c_p[T_2-T_1]_{DBT} \implies -82.39947 =1.005[T_2-31]_{DBT}"

"[T_2]_{DBT}=-50.98952 ^0C"

"\bar{q}= \frac{12.5}{0.1517}=-82.39947" The negative sign shows heat is absorbed.

"q=c_p[T_2-T_1]_{WBT} \implies -82.39947 =1.005[T_2-18.5]_{WBT}"

"[T_2]_{WBT}=--63.48952 ^0C"